OpenStudy (anonymous):

Solve the equation:

OpenStudy (anonymous):

\[x^3-(\sqrt{2}+1)x^2+2=0\]

OpenStudy (blaque678):

I got \[\frac{ \sqrt{2}+i \sqrt{2} }{ 2 } and \frac{ \sqrt{2} -i \sqrt{2}}{ 2 }\]

OpenStudy (anonymous):

Can you show me how you got that?